OpenStudy (anonymous):

hey could someone please explain how normal subgroups work? with some good examples if you dont mind

OpenStudy (anonymous):

Yep, absolutely. A normal subgroup is a subgroup N of a group G such that every element of N, when conjugated by an element of G, goes back to another element of N. Conjugation meaning multiplication of the following type: \(g^{-1}ng\). This is a pretty abstract definition that isn't super easy to understand right away, so some examples are definitely in order. First example: you've probably already heard that every subgroup of an abelian group is normal. This becomes pretty clear when you look at the definition. If you want \(g^{-1}Ng\to N\), and the subgroup is abelian, you can just move the \(g\) through by commuting the multiplication, giving you \(g^{-1}gN\), which cancels to N. Second example: The kernel of a homomorphism is always a normal subgroup (and vice versa, every normal subgroup is the kernel of some homomorphism). Here are a couple of classical examples. One is the kernel of the determinant homomorphism on GLn(R). That kernel is the group SLn(R), which is normal in GLn(R). Another example is the sign homomorphism of Sn. The kernel of that homomorphism is An, which is a normal subgroup of Sn. Hopefully that helps a bit, let me know if you have other specific questions about the concept.

OpenStudy (anonymous):

hmm yeah it does i am trying to do this one practice problem which i dont quite get at all would you mind helping me out with it like yeah i know the definition and your explanation filled in some blanks i had

OpenStudy (anonymous):

so basically if that group is abelian then it has to be normal?

OpenStudy (anonymous):

Yeah, every subgroup of an abelian group is always normal, because you can just commute the multiplication.

OpenStudy (anonymous):

but how would we find out if one is abelian

OpenStudy (anonymous):

Computation, mostly.

OpenStudy (anonymous):

Is there a direct method of finding if a group is abelian besides computing elements?

OpenStudy (anonymous):

In general, if you conjugate in an abelian group, it doesn't do anything. Proof:\[ g,h\in G\text{ (abelian)}\\ gh=hg \text{ by abelian property}\\ ghg^{-1}=hgg^{-1}\text{ multiply on the right by g inverse}\\ ghg^{-1}=h \text{ cancel} \]So, conjugation in an abelian group sends every element to itself. As far as finding out whether a group is abelian, generally you can just look the group up, or think about some examples. If a group is non-abelian it's usually pretty easy to come up with an example or two of elements that don't commute with each other.

OpenStudy (anonymous):

Important examples to know are that Sn for n>2 is always non-abelian, Z/n is always abelian, GLnR and SLnR are non-abelian.

OpenStudy (anonymous):

could we try one example that i got

OpenStudy (anonymous):

Except for when n=1, then GLnR and SLnR are abelian.

OpenStudy (anonymous):

Yeah, go ahead and post it.

OpenStudy (anonymous):

1) The dihedral group of order 12 is D6 = {e, r, r^2, r^3, r^4, r^5, s, rs, (r^2)s, (r^3)s, (r^4)s, (r^5)s } where |r| = 6, |s| = 2 and sr = (r^5)s a) Let H = <r^2> and show that H is normal to D6( the sideway triangle symbol). b) Construct the Caylet table of D6/H. Is this group abelian? Is it cyclic?

OpenStudy (anonymous):

D6 is non-abelian. In general, the dihedral groups are non-abelian. They're abelian for n<3, but those groups aren't usually referred to as dihedral.

OpenStudy (anonymous):

So, with that in mind, to show normal you have to take a different approach. First step, I would compute some conjugations of that group to get an idea for what it looks like.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

Oh, by the way, a good way to know that D6 is non-abelian even if you didn't already know it is that they told you in the question. When they said \(sr=r^5s\), that means it's non-abelian because \(sr\ne rs\).

OpenStudy (anonymous):

oo yeah thats truee but how do we show that H is normal to D6? well clearly it is nt because it is non abelian

OpenStudy (anonymous):

is that sufficient enough for 1a ? :s

OpenStudy (anonymous):

H is normal to D6. Abelian implies normal, but the converse isn't true. So, what you need to do is try conjugating elements of H by elements of D6, and see what happens.

OpenStudy (anonymous):

what do you mean by that could we do one element from D6? and try it

OpenStudy (anonymous):

Yeah, take elements from D6 and conjugate H by them, and see what you get as a result

OpenStudy (anonymous):

Start with r as an example. Conjugate each element of H by r.

OpenStudy (anonymous):

im sorry @nbouscal but i still dont get it

OpenStudy (anonymous):

Okay, well, first off, can you explicitly list the elements of H?

OpenStudy (anonymous):

i have done using the left and right cosets with permutation groups but not with like r and r^2 and stuff

OpenStudy (anonymous):

H = (r^2)

OpenStudy (anonymous):

H is the subgroup generated by r^2, so it's more than just that one element.

OpenStudy (anonymous):

Remember what is required to be a subgroup: it must be closed under the multiplication and closed under inverses, and it must contain the identity.

OpenStudy (anonymous):

alright so what i think it is eH rH r^2H .... ?

OpenStudy (anonymous):

eH = (r^2) rH = (r^3) r^2H=(r^4) ? and i do this for all the elements or am i totally wrong

OpenStudy (anonymous):

No, not at all that. We're still just looking for what the elements of H are to begin with.

OpenStudy (anonymous):

H is a subgroup, so it has to fulfill all the requirements to be a subgroup. It's generated by r^2, so that means it has all powers of r^2 in it.

OpenStudy (anonymous):

First step is it has to have e, because every subgroup has to have the identity. Then we can also look and see, what is the inverse of r^2? Because it has to be closed under inverses.

OpenStudy (anonymous):

but doesnt for one group to be normal the left coset has to equal to the right coset?

OpenStudy (anonymous):

Yeah, we're not that far yet though, we still have to find the subgroup itself to begin with

OpenStudy (anonymous):

OO are we showing first if it is a group by showing associativity inverse and identity?

OpenStudy (anonymous):

We're just finding out what the group is. What are the elements of H?

OpenStudy (anonymous):

We have to be able to list out the elements of H explicitly, so that we know what we're multiplying by when we figure out cosets

OpenStudy (anonymous):

what do you mean by writing it explicitly i dont get that

OpenStudy (anonymous):

I mean, what are the elements of the group? We have listed the 12 elements of D6, we need to list the elements of H.

OpenStudy (anonymous):

but the element of H is given to us which is (r^2) is it not

OpenStudy (anonymous):

No. If H was just r^2, it wouldn't be a subgroup! Subgroups have to have the identity, and have to be closed under multiplication and inverses. Just {r^2} is not a subgroup! What was given to us is <r^2>, which means the subgroup generated by r^2.

OpenStudy (anonymous):

Perhaps you ought to review what generator terminology means, what it means to have a group generated by an element. Cyclic subgroups and cyclic groups.

OpenStudy (anonymous):

gosh this is confusing i will try learning more of this stuff and post questions later on then

OpenStudy (anonymous):

Okay. To answer the question for you, H is {e, r^2, r^4}. Definitely review generators, subgroups, cyclic subgroups, etc.

OpenStudy (anonymous):

@hamza_b23, did my response to your question about injective, surjective, and bijective help? Just curious.

OpenStudy (anonymous):

yeah it did help a lot thank you